# Design Space Exploration of Embedded SoC Architectures for Real-Time Optimal Control

### v1

@@ IV-B Programming Interface and Libraries
In order to create a unified platform for comparing the performance of differing architectures for Robotic/DSP applications, we created a C library of commonly used operators, matlib [^1] https://github.com/kris-dong/matlib. Matlib provides a lightweight interface to various linear algebra operations similar to Eigen [10], a C++ header-only library. matlib was used to write reference implementations for each accelerator backend; however, to achieve full performance, hand-tuned implementations required optimizations across the abstractions provided by these function definitions.
@@ V Hardware Software Co-optimization
Ensuring each architecture is optimized to the extreme to showcase the potential performance benefits and limitations inherent in each architectural choice. Both architectural development and software optimization techniques are deployed onto the configurable backends, i.e. vector cores Saturn and systolic arrays Gemmini. These optimizations are essential not only for a fair comparison between different hardware architectures but also for isolating architectural characteristics without error from sub-optimal software implementations. By pushing the software to its performance limits, we eliminate variables, such as inefficient software mappings or overheads introduced by inefficiently written code, that might obscure the hardware’s true capabilities. This approach ensures that any observed performance differences are directly attributable to the hardware’s architectural design rather than extrinsic software factors.
@@ V-A Optimizations on Vector Core Saturn
To initially accelerate TinyMPC, we utilized a library-based approach. For every matlib function that was used within the solver, we wrote a vectorized implementation using RVV intrinsic support from gcc. However, as shown in Figure 3, although the vectorized matlib code had meaningful speedup over scalar matlib code running on Rocket, highly optimized scalar code using Eigen still outperformed the Saturn implementation, necessitating further optimizations.
@@ V-A1 Software Loop Unrolling
One benefit of the RVV ISA is the ability to perform register grouping using the LMUL field. This performs register grouping and unrolls contiguous vector operations in hardware using a vector sequencer, theoretically eliminating the need for software loop unrolling. As depicted in Figure 4, utilizing LMUL improved the performance of strip-mining operations in matlib. However, increasing LMUL actually harmed performance for the iterative components of TinyMPC, backward_pass and forward_pass. The major component was that these operations had to compute GEMV and vector addition operations serially and could not be mapped to larger vector registers due to dependencies. Furthermore, optimized GEMV kernels on Saturn require using the vfmacc_vf instruction, multiplying a scalar element of a vector against a column of a matrix. However, due to TinyMPC using vector dimensions of 4 and 12 for the iterative kernels, these could only be partially mapped to the 512-bit vector registers, further hindering the use of LMUL. Aggressive software loop unrolling allowed variation in the scalar elements across instructions, compensating for the absence of hardware unrolling.
Alternatively, we considered using vector summation using reduction operators (vfred(o|u)sum) to increase the utilization of vector lanes for wide matrices. However, Saturn currently implements vector reduction serially. Because of this, despite using fewer vector lanes, vfmacc_vf is better suited than vector reduction.
@@ V-B1 Build a Systolic Array that supports GEMV
We architect Gemmini’s systolic array to support both GEMM and GEMV operations and effectively address the previously limited utilization of processing elements. This expansion of the systolic array change ensures a fair comparison between Saturn and Gemmini. The rationale is that Saturn, as a general proposed-vector machine, supports general matrix-vector multiplication(GEMV) and general matrix-matrix multiplication (GEMM) well. However, Gemmini is facing the limitation of only 1/DIM utilization in GEMV operations since only one column of the mesh PEs will be utilized. For example, the utilization is approximately 25% for a mesh with a dimension of 4x4. Systolic array’s utilization decreases when the mesh size increases.
The primary difficulty with the matrix-vector operation $AB+D$ is the lack of reuse with the $A$ matrix. No longer can we propagate $A$ rightwards, as each processing element (PE) requires a new element every cycle. Given the desire for full utilization of the mesh, we first added additional scratchpad banks and modified the execute controller to fetch $DIM\times DIM$ elements of $A$ in parallel. By providing mesh $DIM^{2}$ elements of $A$ at every compute cycle, one for each of the PE in the systolic array. Next, to improve the reuse of B, we broadcast a singular element of B across all columns. This allows the columns to be independent of one another during GEMV operations. The changes allow all the PEs to be utilized to compute $DIM^{2}$ of A ×$\times DIM$ elements of $B$ for GEMV operation in the systolic array 7.
By re-architecting the original Gemmini, the modified GEMV Gemmini achieves approximately more than 4x speedup in GEMV by achieving full utilization on such operation while maintaining its’ performance in GEMM. Another benefit is removing the delay across columns, as each input can be directly wired to a PE with a mux. The trade-off is that the additional features come at a 2% area cost in a 4x4 mesh compared to Gemmini’s original design. As depicted in Table II, an additional area is attributed to the extra scratchpad banks since our feature needs at least $DIM+1$ banks to ensure all the necessary data for the mesh being loaded in one cycle. $DIM$ for the elements of input $A$ and $1$ for the biases in weight stationary and weights(vector) $B$ in output stationary.
Figure 6: Example $DIM=4$ scratchpad setup with GEMV mesh. The rows of A are strided across $DIM$ banks to allow for $DIM\times DIM$ elements to be accessed at once. The weight vector is stored consecutively in memory.
@@
The heatmap in Figure 8 illustrates the speedup achieved by our design compared to the original Gemmini design over a range of matrix and vector sizes. The x-axis represents the matrix width and the vector length (K), while the y-axis shows the matrix height (I). Each cell depicts the relative speedup obtained for a specific combination of matrix and vector dimensions. The result demonstrates that the new GEMV/GEMM systolic array achieves a significant performance improvement over the traditional GEMM systolic array on various size of vector and matrix sizes. On average, the new architecture delivers on average ~6x speedup on GEMV operations while maintaining similar performance for General Matrix-Matrix Multiplication (GEMM).
@@ V-B2 Perform Software Optimization on Gemmini
Static mapping of kernels involves pre-calculating and optimizing the allocation of data and tasks to various processing elements within the hardware. In scenarios involving fixed-size operations like those commonly found in Model Predictive Control (MPC), dimensions, tiling, and indexing can be determined without dynamic memory allocation. This allows for the entire computational workflow to be streamlined, as all necessary instructions—mvin (memory vector in), execute, and mvout (memory vector out)—can be prepared and optimized beforehand. This approach not only simplifies the coding and execution process but also significantly reduces the computational overhead associated with dynamically calculating these parameters during runtime.
@@
Beyond improving throughput, this method also involves the removal of redundant operations that do not contribute to the result of the computation. This includes unnecessary configuration commands, fencing operations that ensure data integrity but may be excessive, and redundant data moves (mvin) of data already present in the system. By eliminating these operations, the efficiency of the execution process is further enhanced. Each instruction executed is necessary and directly contributes to the computational goal, minimizing the cycle time and improving the overall speed of computation.
@@ V-B3 Use of Gemmini’s Fine-Grained ISA
Gemmini’s coarse-grained instruction set is necessary for most machine learning applications to achieve high mesh utilization. However, coarse-grained instructions require multiple RoCC instructions for configuration, spending 5-7 instructions before beginning execution. Additionally, constructing these RoCC instructions can take multiple cycles due to the bit-shifting that needs to be performed by the scalar CPU to construct the RoCC arguments. For the matrix sizes found in TinyMPC, coarse-grained instructions only map to 2-6 fine-grained instructions, resulting in negligible benefit from hardware instruction sequencing. Furthermore, coarse-grained instructions require the input operands to be stored in memory, preventing additional optimizations from re-using data on Gemmini’s scratchpad.
However, using the fine-grained ISA requires improved instruction throughput from the CPU, driving the RoCC interface to achieve high utilization of Gemmini. To address this, we aggressively unroll code, and perform indexing and address calculation at compile-time using static mapping to minimize the overhead of constructing RoCC instructions.
@@ V-B4 Scratchpad-Resident Linear Algebra Operations
In order to support data-dependent operations on Gemmini, modifying the typical DNN kernel mapping for Gemmini was necessary. Typically, inputs to an operation are loaded into Gemmini’s scratchpad before they are used as inputs to tiled operations, accumulating in Gemmini’s accumulator memory. These results are then stored back in DRAM before being reused in subsequent operations. However, when performing operations that takes several cycles on Gemmini, this overhead severely limits performance. In addition to the latency induced by data movement, explicit fence instructions must be inserted between Gemmini store and load instructions, as Gemmini’s ROB does not track data RAW hazards across memory operations. In our experimentation, this can introduce up to 600 cycles of stalling upon a fence instruction within the iterative phase of TinyMPC.
@@
To avoid this issue, we perform several optimizations to our mapping. First, we load all matrices used by TinyMPC onto the first bank of the scratchpad, in addition to several utility matrices, such as identity and negative identity matrices. Secondly, intermediate results computed by Gemmini are directly written to the scratchpad so that they can be immediately reused in subsequent operations. One downside of this approach is that writing to the scratchpad instead of the main memory prevents the use of Gemmini’s output scaling pipeline, which can be helpful in performing fused scaling/GEMM operations. In order to compensate for this, we allocate additional utility matrices onto the scratchpad, which are commonly used scalar multiples of the identity matrix.
@@ V-B6 Activation Functions and Pooling
Finally, computing the residuals of TinyMPC requires calculating a global maximum across vectors of approximately 100 elements each. While this could potentially implemented using sequences of ReLU operations, computing a maximum between two arbitrary vectors would require over 5 operations, achieving comparable performance to simply using a scalar CPU. However, to avoid computing the entire maximum on the scalar core, Gemmini’s hardware support for max-pooling and be utilized. By using a pool-size of 2 when moving out to shared memory, Gemmini can perform a reduction across 4 scratchpad rows. This can potentially be increased by increasing the hardware pooling dimensions. Although this does not enable max reduction within a scratchpad row, this still reduces the max reduction needed to be performed on the CPU by a factor of 4.
@@ VI Evaluation
We conduct a detailed profiling and synthesis to explore the trade-offs between area, performance, and utilization of each backend. First, we carried out the comparison over standard benchmarks between RISC-V cores, vector core Saturn, and the systolic array, Gemmini, specifically General Matrix-Vector Multiplication (GEMV) and General Matrix-Matrix Multiplication (GEMM). Second, we particularly evaluated each backend in the context of robotics applications with the target workloads. By synthesizing different architectural configurations, we understand the strengths and limitations of each approach under varying computational loads, and analyze the trade-off between performance, area, and utilization across all designs.
@@ VI-A1 Kernel-level Evaluation
The heatmap in Figure 13 and Figure 15 shows the performance patterns of vector architecture and systolic array on different operations. Each backend has its unique strength over the other with different operations.
Figure 13 shows the speedup achieved by Saturn compared to the original Gemmini design when performing GEMV operations over a range of randomly generated matrix and vector sizes. Here both backends are driven by a simple in-order RISC-V core, Rocket[1]. The x-axis represents the matrix width and the vector length(K), while the y-axis shows the matrix height (I). Each cell depicts the relative speedup obtained for a specific combination of matrix and vector dimensions. On average, Saturn delivers on average ~2.78x speedup on GEMV operations, as Gemmini’s original architecture can only utilize one column of its mesh for this workload.
After adding the GEMV support, Gemmini’s performance on GEMV operations improved significantly by over 5x on average, which allows Gemmini’s to outperform Saturn by 2.34x in various vector and matrix sizes. However, Saturn must perform matrix-vector operations using vector-scalar operations, broadcasting one element of the vector to perform parallel reductions. The GEMV kernel uses register grouping with LMUL=8 to reduce instruction throughput requirements from the front end. However, driven by a simple Rocket core, for small matrix heights (I), this kernel cannot make full use of the grouped registers. In this case, Gemmini, which can perform sequenced operations using the K dimension, performs comparably to Saturn.
@@
The result demonstrates that the distinct strength of both types of architecture. In general, the vector core performs better than the original Gemmini for GEMV operations, and the systolic array achieves a significant performance for GEMM since it exploits the 2D spacial attribute of matmul operation.
@@ VI-A2 End-to-end Workload Evaluation
Breaking down the performance within each category, we profiled representative architecture candidates from each categories17: RISC-V CPU cores, RISC-V CPU with vector extensions Saturn[28], and the CISC-style systolic array Gemmini[9], where RISC-V CPU cores includes simple in-order core [1], superscalar in-order core, out-of-order BOOM core[3], a more performant out-of-order MegaBOOM core with 2FPUs[29]. When comparing end-to-end performance of control workloads, we use Rocket running optimized sclaar Eigen code as a baseline. However, when evaluating the performance of individual kernels, we use the matlib implementation of the kernel; although matlib is not as performant as Eigen, our matlib implementation has the same interface as the accelerated implementation of each kernel.
The different BOOM configurations showcase varying specifications designed for specific performance needs. The Small BOOM has a fetch width of 4, a decode width of 1, and three instruction queues (IQ) for memory (mem), integer (int), and floating-point (fp) pipelines. The Mid BOOM also features a fetch width of 4 but increases the decode width to 2, enhancing its IQ configurations with mem having 1 issue and 2 dispatches, int with 2 issues and 2 dispatches, and fp with 1 issue and 2 dispatches. The Large BOOM continues with a fetch width of 4 and a decode width of 1, increasing dispatch capacity with mem, int, and fp pipelines featuring more dispatch capabilities. The Mega BOOM expands further with a fetch width of 8 and a decode width of 4, having each pipeline’s IQ equipped for higher issue and dispatch rates.
@@
In operations where high data parallelism is exploitable, Saturn tends to perform better, which is evident from the consistent high speedup factors, particularly in the last chart, where it achieves a very high speedup in several operations. In these cases, Saturn can fully map operations to vector registers without significant overhead from handling tail-cases and can make use of instruction sequencing and register grouping.
@@
Gemmini’s performance peaks in specific operations like forward passes and linear cost updates, as shown in chart18, suggesting that its architecture is highly optimized for scenarios where matrix-vector operations are dominant. The significant variability in Gemmini’s performance across different operations may indicate its specialized nature, excelling greatly in its niche (matrix operations) but perhaps not as flexible or efficient for other types of data processing tasks compared to general-purpose vector processors like Saturn.
@@
Furthermore, we conducted an initial comparison between Gemmini and Saturn to better understand each architecture’s relative performance and their unique strengths in certain types of operations. The configurations are carefully chosen, such as Gemmini with 4x4 FP Mesh and Saturn V512D512, both driven by Rocket. These configurations ensure the backends of each architecture have equal amounts of PEs, thus allowing them to perform theoretically the same FLOPs/cycle. Besides the same computational capacity of the backend of each candidate, choosing the front end to be Rocket for both also allows the instructional throughput to be relatively similar among both candidates.
@@
Due to the flexible nature of the RVV vector instruction set as well as the Saturn generator, we were able to evaluate many hardware configurations using the same software mapping. By using RVV intrinsics and dynamically computing VLMAX, the same binary can be executed on ever hardware configuration evaluated. Due to the fact that the iterative functions within TinyMPC can often only utilize 4 vector lanes at most, Saturn configurations with a DLEN of 128 are more efficient compared to their counterparts with a DLEN of 256. Furthermore, achieving the highest optimal performance required both a superscalar Shuttle frontend and a DLEN=256 backend. However, interestingly, the single-issue reference design outperformed the DSP and general-purpose Saturn configurations for a DLEN=256, despite the fact that both of those configurations should outperform the single-issue reference design. Further investigations could include evaluating more area-minimal Saturn configurations. Currently, under 1.4mm2, a Rocket core is the most efficient implementation. However, minimal Saturn configurations could result in improved performance in this domain due to Saturn’s instruction sequencing. For higher-performance configurations, hardware support for explicitly operating on small, dense matrices could improve Saturn’s performance as the vector and datapath lengths scale beyond what was evaluated in this work.
Evaluating multiple configurations for Gemmini was more challenging, as programming Gemmini using its low-level assembly interface requires manual rewriting across changes to the mesh. Because of this, we primarily focused on creating a single, well-optimized implementation for the OSGemminiRocket configuration, which uses a 4$\times$4 output-stationary FP32 mesh, evaluated with both a 64KB and 32KB scratchpad. Furthermore, we include the area report for an equivalent weight stationary design, which requires the generation of a 1KB accumulator memory for computing intermediate results. However, this implementation has significantly worse performance due to the fact that the software optimizations aside from software unrolling and static mapping have not been implemented for this design. Under an area window of 1.5mm2 to $2^{3}$, Gemmini is the optimal design, even without the use of hardware support for GEMV operations. In this case, the efficiency improvements could be due to the ability to fuse scaling and matrix operations, alongside the fact that the fine-grained instruction sequencing works well with blocks of matrices that are multiples of 4. To achieve a more thorough understanding of Gemmini’s end-to-end performance characteristics, future evaluations can consider other mesh dimensions, smaller scratchpad capacities, as well as hardware modifications such as the GEMV support presented in this work, as well as activation and pooling functions suitable for classical control workloads.

### v2

@@ IV-B Programming Interface and Libraries
To create a unified platform for comparing the performance of differing architectures for Robotic and DSP applications, we create a C library of commonly used operators, matlib [^1] https://github.com/ucb-bar/matlib. matlib provides a lightweight interface to various linear algebra operations similar to Eigen [10], a C++ header-only library. matlib was used to write reference implementations for each backend; however, to achieve full performance, hand-tuned implementations required optimizations across the abstractions provided by these function definitions.
@@ V Hardware Software Co-optimization
Developing highly optimized software mapping for each hardware target is necessary. These optimizations ensure a clear comparison between different hardware architectures by isolating architectural characteristics from sub-optimal software implementations. This approach attempts to overcome the impact of other variables such as inefficient software mappings or overheads introduced by inefficiently written code, that might obscure the hardware’s true capabilities. This approach ensures that any observed performance differences are directly attributable to the hardware’s architectural design rather than extrinsic software factors.
@@ V-A Optimizations on Vector Core Saturn
To initially accelerate TinyMPC, we utilize a library-based approach. For every matlib function that is used within the solver, we write a vectorized implementation using RISC-V vector extension (RVV) intrinsic support from gcc. However, as shown in Figure 3, although the vectorized matlib code has meaningful speedup over scalar matlib code running on Rocket, highly optimized scalar code using Eigen still outperforms the Saturn implementation, necessitating further optimizations.
@@ V-A1 Software Loop Unrolling
One benefit of the RVV ISA is its ability to perform register grouping using the LMUL field. This performs register grouping and unrolls contiguous vector operations in hardware using a vector sequencer, theoretically eliminating the need for software loop unrolling. As depicted in Figure 4, utilizing LMUL improved the performance of strip-mining operations in matlib. However, increasing LMUL harmed performance for the iterative components of TinyMPC, backward_pass and forward_pass. The major component was that these operations must to compute GEMV and vector addition operations serially and could not be mapped to larger vector registers due to dependencies. Furthermore, optimized GEMV kernels on Saturn require using the vfmacc_vf instruction, multiplying a scalar element of a vector against a column of a matrix. However, due to TinyMPC using vector dimensions of 4 and 12 for the iterative kernels, these could only be partially mapped to the 512-bit vector registers, further hindering the use of LMUL. Aggressive software loop unrolling allowed variation in the scalar elements across instructions, compensating for the absence of hardware unrolling.
Alternatively, we consider using vector summation using reduction operators (vfred(o|u)sum) to increase the utilization of vector lanes for wide matrices. However, Saturn currently implements vector reduction serially. Because of this, despite using fewer vector lanes, vfmacc_vf is better suited than vector reduction.
@@ V-B1 Build a Systolic Array that supports GEMV
We architect Gemmini’s systolic array to support both GEMM and GEMV operations and effectively address the previously limited utilization of processing elements. This expansion of the systolic array change ensures a fair comparison between Saturn and Gemmini. The rationale is that Saturn, as a general proposed-vector machine, supports GEMV and GEMM well. However, Gemmini is facing the limitation of only 1/DIM utilization in GEMV operations since only one column of the mesh PEs will be utilized. For example, the utilization is approximately 25% for a mesh with a dimension of 4x4. The utilization of the systolic array decreases when the mesh size increases.
[…]
By re-architecting Gemmini, the modified GEMV Gemmini achieves more than a 4$\times$ speedup in GEMV by achieving full utilization on such operation while maintaining its performance in GEMM. Another benefit is removing the delay across columns, as each input can be directly wired to a PE with a multiplexer. The trade-off is that the additional features come at a 2% area cost in a 4$\times$4 mesh compared to Gemmini’s original design. As depicted in Table II, an additional area is attributed to the extra scratchpad banks since our feature needs at least $\textrm{DIM}+1$ banks to ensure all the necessary data for the mesh being loaded in one cycle. DIM for the elements of input $A$ and $1$ for the biases in weight stationary mode and weights (vector) $B$ in output stationary mode.
[…]
@@
The heatmap in Figure 8 illustrates the speedup achieved by our design compared to the original Gemmini design over a range of matrix and vector sizes. The x-axis represents the matrix width and the vector length ($K$), while the y-axis shows the matrix height ($I$). Each cell shows the speedup achieved for the given matrix and vector dimensions. The result demonstrates that the new GEMV/GEMM systolic array achieves a significant performance improvement over the traditional GEMM systolic array on various sizes of vector and matrix sizes. On average, the new architecture delivers on average ~a 6$\times$ speedup on GEMV operations while maintaining similar performance for GEMM.
@@ V-B2 Perform Software Optimization on Gemmini
Static mapping of kernels involves pre-calculating and optimizing the allocation of data and tasks to various processing elements within the hardware. In scenarios involving fixed-size operations like those commonly found in MPC, dimensions, tiling, and indexing can be determined without dynamic memory allocation. This allows for the entire computational workflow to be streamlined, as all necessary instructions such as moving data in (i.e. mvin) and out (i.e. mvout) of scratchpad and computation (i.e. execute) can be prepared and optimized beforehand. This approach not only simplifies the coding and execution process but also significantly reduces the computational overhead associated with dynamically calculating these parameters during runtime.
@@
Beyond improving throughput, this method also involves the removal of redundant operations that do not contribute to the result of the computation. This includes unnecessary configuration commands, fencing operations that ensure data integrity but may be excessive, and redundant data movement of data already present in the system. By eliminating these operations, the efficiency of the execution process is further enhanced. Each instruction executed is necessary and directly contributes to the computational goal, minimizing the cycle time and improving the overall speed of computation.
@@ V-B3 Use of Gemmini’s Fine-Grained ISA
Gemmini’s coarse-grained instruction set is necessary for most machine-learning applications to achieve high mesh utilization. However, coarse-grained instructions require multiple RoCC instructions for configuration, spending 5-7 instructions before beginning execution. Additionally, constructing these RoCC instructions can take multiple cycles due to the bit-shifting that needs to be performed by the scalar CPU to construct the RoCC arguments. For the matrix sizes found in TinyMPC, coarse-grained instructions only map to 2-6 fine-grained instructions, resulting in negligible benefit from hardware instruction sequencing. Furthermore, coarse-grained instructions require the input operands to be stored in memory, preventing additional optimizations from re-using data on Gemmini’s scratchpad.
However, using the fine-grained ISA requires improved instruction throughput from the CPU, driving the RoCC interface to achieve high utilization of Gemmini. To address this, we aggressively unroll code and perform indexing and address calculation at compile-time using static mapping to minimize the overhead of constructing RoCC instructions.
@@ V-B4 Scratchpad-Resident Linear Algebra Operations
In order to support data-dependent operations on Gemmini, modifying the typical DNN kernel mapping for Gemmini was necessary. Typically, inputs to an operation are loaded into Gemmini’s scratchpad before they are used as inputs to tiled operations, accumulating in Gemmini’s accumulator memory. These results are then stored back in DRAM before being reused in subsequent operations. However, when performing operations that take several cycles on Gemmini, this overhead severely limits performance. In addition to the latency induced by data movement, explicit fence instructions must be inserted between Gemmini store and load instructions, as Gemmini’s ROB does not track data RAW hazards across memory operations. In our experimentation, this can introduce up to 600 cycles of stalling upon a fence instruction within the iterative phase of TinyMPC.
@@
To avoid this issue, we perform several optimizations to our mapping. First, we load all matrices used by TinyMPC onto the first bank of the scratchpad, in addition to several utility matrices, such as identity and negative identity matrices. Secondly, intermediate results computed by Gemmini are directly written to the scratchpad so that they can be immediately reused in subsequent operations. One downside of this approach is that writing to the scratchpad instead of the main memory prevents the use of Gemmini’s output scaling pipeline, which can help perform fused scaling/GEMM operations. To compensate for this, we allocate additional utility matrices onto the scratchpad, which are commonly used scalar multiples of the identity matrix.
@@ V-B6 Activation Functions and Pooling
Finally, computing the residuals of TinyMPC requires calculating a global maximum across vectors of approximately 100 elements each. While this could potentially implemented using sequences of ReLU operations, computing a maximum between two arbitrary vectors would require over 5 operations, achieving comparable performance to simply using a scalar CPU. However, to avoid computing the entire maximum on the scalar core, Gemmini’s hardware support for max-pooling and be utilized. By using a pool size of 2 when moving out to shared memory, Gemmini can perform a reduction across 4 scratchpad rows. This can potentially be increased by increasing the hardware pooling dimensions. Although this does not enable max reduction within a scratchpad row, this still reduces the max reduction needed to be performed on the CPU by a factor of 4.
@@ VI Evaluation
We conduct a detailed profiling and synthesis to explore the trade-offs between area, performance, and utilization of each backend. First, we carried out the comparison over standard benchmarks between RISC-V cores, vector core Saturn, and the systolic array, Gemmini, specifically GEMV and GEMM operations. Second, we evaluated each backend in the context of robotics applications with the target workloads. By synthesizing different architectural configurations, we understand the strengths and limitations of each approach under varying computational loads and analyze the trade-off between performance, area, and utilization across all designs.
@@ VI-A1 Kernel-level Evaluation
The heatmaps in Figure 13 and Figure 15 illustrate the performance patterns of the vector architecture and systolic array across various operations, highlighting each backend’s strengths. These differences in performance reflect that each backend has its unique strength over the other with different operations.
@@
Figure 13 demonstrates the speedup achieved by Saturn over the original Gemmini design for GEMV operations across a range of randomly generated matrix and vector sizes. Both backends are powered by a simple in-order RISC-V core, Rocket[1]. The x-axis represents the matrix width and vector length ($K$), while the y-axis indicates the matrix height ($I$). Each cell displays the relative speedup achieved for a given matrix-vector size combination. On average, Saturn achieves a speedup of approximately 2.78$\times$ for GEMV operations, as the original Gemmini architecture can only utilize a single column of its mesh for this workload.
After adding the GEMV support, Gemmini’s performance on GEMV operations improved significantly by over 5$\times$ on average, which allows Gemmini’s to outperform Saturn by 2.34$\times$ in various vector and matrix sizes. However, Saturn must perform matrix-vector operations using vector-scalar operations, broadcasting one element of the vector to perform parallel reductions. The GEMV kernel uses register grouping with LMUL=8 to reduce instruction throughput requirements from the front end. However, driven by a simple Rocket core, for small matrix heights ($I$), this kernel cannot make full use of the grouped registers. In this case, Gemmini, which can perform sequenced operations using the K dimension, performs comparably to Saturn.
@@
The result demonstrates the distinct strength of both types of architecture. In general, the vector core performs better than the original Gemmini for GEMV operations, and the systolic array achieves a significant performance for GEMM since it exploits the 2D spatial attribute of the GEMM operation.
@@ VI-A2 End-to-end Workload Evaluation
Breaking down the performance within each category, we profiled representative architecture candidates from each category: RISC-V CPU cores, RISC-V CPU with vector extensions Saturn, and the systolic array Gemmini, where RISC-V CPU cores include simple in-order core, superscalar in-order core, out-of-order BOOM core, a more performant out-of-order MegaBOOM core with 2FPUs. When comparing the end-to-end performance of control workloads, we use Rocket running optimized scalar Eigen code as a baseline. However, when evaluating the performance of individual kernels, we use the matlib implementation of the kernel; although matlib is not as performant as Eigen, our matlib implementation has the same interface as the accelerated implementation of each kernel.
The different BOOM configurations showcase varying specifications designed for specific performance needs. The Small BOOM has a fetch width of 4, a decode width of 1, and three instruction queues (IQ) for memory (mem), integer (int), and floating-point (fp) pipelines. The Medium BOOM also features a fetch width of 4 but increases the decode width to 2, enhancing its IQ configurations with mem having 1 issue and 2 dispatches, int with 2 issues and 2 dispatches, and fp with 1 issue and 2 dispatches. The Large BOOM continues with a fetch width of 4 and a decode width of 1, increasing dispatch capacity with mem, int, and fp pipelines featuring more dispatch capabilities. The Mega BOOM expands further with a fetch width of 8 and a decode width of 4, having each pipeline’s IQ equipped for higher issue and dispatch rates.
@@
In operations where high data parallelism is exploitable, Saturn tends to perform better, which is evident from the consistent high speedup factors, particularly in the last chart, where it achieves a very high speedup in several operations. In these cases, Saturn can fully map operations to vector registers without significant overhead from handling tail cases and can make use of instruction sequencing and register grouping.
@@
Gemmini’s performance peaks in specific operations like forward passes and linear cost updates, as shown in Figure 18, suggesting that its architecture is highly optimized for scenarios where matrix-vector operations are dominant. The significant variability in Gemmini’s performance across different operations may indicate its specialized nature, excelling greatly in its niche (matrix operations) but perhaps not as flexible or efficient for other types of data processing tasks compared to general-purpose vector processors like Saturn.
@@
Furthermore, we conduct an initial comparison between Gemmini and Saturn to better understand each architecture’s relative performance and their unique strengths in certain types of operations. The configurations are carefully chosen, such as Gemmini with 4$\times$4 FP Mesh and Saturn V512D512, both driven by Rocket. These configurations ensure the backends of each architecture have equal amounts of PEs, thus allowing them to perform theoretically the same FLOPs/cycle. Besides the same computational capacity of the backend of each candidate, choosing the front end to be Rocket for both also allows the instructional throughput to be relatively similar among both candidates.
@@
Due to the flexible nature of the RVV vector instruction set as well as the Saturn generator, we were able to evaluate many hardware configurations using the same software mapping. By using RVV intrinsics and dynamically computing VLMAX, the same binary can be executed on every hardware configuration evaluated. Since the iterative functions within TinyMPC can often only utilize 4 vector lanes at most, Saturn configurations with a DLEN of 128 are more efficient compared to their counterparts with a DLEN of 256. Furthermore, achieving the highest optimal performance required both a superscalar Shuttle frontend and a DLEN=256 backend. However, interestingly, the single-issue reference design outperformed the DSP and general-purpose Saturn configurations for a DLEN=256, even though both of those configurations should outperform the single-issue reference design. Further investigations could include evaluating more area-minimal Saturn configurations. Currently, under 1.4mm2, a Rocket core is the most efficient implementation. However, minimal Saturn configurations could result in improved performance in this domain due to Saturn’s instruction sequencing. For higher-performance configurations, hardware support for explicitly operating on small, dense matrices could improve Saturn’s performance as the vector and datapath lengths scale beyond what was evaluated in this work.
Evaluating multiple configurations for Gemmini was more challenging, as programming Gemmini using its low-level assembly interface requires manual rewriting across changes to the mesh. Because of this, we primarily focused on creating a single, well-optimized implementation for the OSGemminiRocket configuration, which uses a 4$\times$4 output-stationary FP32 mesh, evaluated with both a 64KB and 32KB scratchpad. Furthermore, we include the area report for an equivalent weight stationary design, which requires the generation of a 1KB accumulator memory for computing intermediate results. However, this implementation has significantly worse performance since the software optimizations aside from software unrolling and static mapping have not been implemented for this design. Under an area window of 1.5mm2 to 2.3mm2, Gemmini is the optimal design, even without the use of hardware support for GEMV operations. In this case, the efficiency improvements could be due to the ability to fuse scaling and matrix operations, alongside the fact that the fine-grained instruction sequencing works well with blocks of matrices that are multiples of 4. To achieve a more thorough understanding of Gemmini’s end-to-end performance characteristics, future evaluations can consider other mesh dimensions, smaller scratchpad capacities, as well as hardware modifications such as the GEMV support presented in this work, as well as activation and pooling functions suitable for classical control workloads.